OpenStudy (anonymous):

if the sides of a square are lengthened by 8 cm, the area becomes 256 cm^2, find the legth of the origional side

OpenStudy (mathteacher1729):

A square has four equal sides, and if each side is of length x, that means the area is x^2. Now, you make each side 8cm longer, so that means each side of the square is now (x+8) and the area is (x+8)^2 or (x+8)(x+8). Since the new area is 256 that means.... (x+8)(x+8)=256 Expand the expression on the left to reveal a quadratic equation, and solve. :)

OpenStudy (anonymous):

12?

OpenStudy (mathteacher1729):

No. If the original side were 12, that means the new sides are 20. 20*20 = 400 . That's too big.

OpenStudy (mathteacher1729):

Question -- "what number , when you square it, is 256?"

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

thats even bigger

OpenStudy (anonymous):

8; the new area is 256, take the square root to get 16 for new length of each side. Since the new side is 16 and the old side was 8 less, the old square was 8 [16-8=8)